# Eye Gaze as a Signal for Conveying User Attention in Contextual AI Systems

### v2

@@ 3.1.3. Protocol
3.2. Contextualizing vision-language model queries with eye tracking information
To begin to explore the value that ET signals can provide in contextual AI systems, we model experiments which reflect potential end-to-end contextual AI systems. In these experiments, we build up historical context by creating timelines of past fixations on physical objects. We include this context in VLM queries and measure the positive impacts on model understanding. Our baseline comparison is a VLM query which uses only the egocentric image as added context (similar to a Ray-Ban Meta or Google Ask Photos query).
We use the Meta Llama 3.2 90B VLM [^3] https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/ (Grattafiori et al., 2024) as our contextual AI agent. We add context to queries with an egocentric image and additional prompting. In both experiments, the agent is constrained via JSON to respond with one option from all currently visible objects, as we are operating under the pretense that in a full system, an object recognition / scene understanding model would be available. Note that this VLM is not fine-tuned, and that a model tuned for egocentric image understanding and / or for a specific task would likely see improved results. Yet, these experiments indicate the added value when incorporating ET contextual information.
3.2.1. E1: “What am I looking at?” with historical context.
3.2.2. E2: ”What am I going to interact with?” with historical context.
We query the VLM “what am I going to interact with?” while again varying the amount of gaze context. In this experiment, we supply the image from a current fixation, where a physical interaction is guaranteed to occur within the next second and at least 10 prior fixations exist (237 distinct image frames). The vast majority of physical interactions in ADT are grasping, pushing, and pulling with hands. This task models the use of ET as supporting context for user action understanding and prediction.
4. Results
4.1. Eye tracking accuracy requirements
Figure 3. Eye tracking accuracy requirements to place gaze accurately within in the ADT dataset, where users performed household actions in an indoor environment (Pan et al., 2023) Near-field and mid-field measurements consider all objects present in the user’s field of view, where interacted objects are being actively manipulated by the user, and fixated objects consider those being directly gazed at.
The near-field ($\leq$ 1 meter) and mid-field (1 - 2 meters) measurements reflect the visual FOV occupied by every object in the environment within the distance threshold. Alternatively, the interacted measurement only aggregates objects when being manually interacted with, including picking up, pushing, pressing, etc. Fixation measurements consider the objects within 2 meters at the time of fixation. Considering ADT’s household scenarios, the interacted and fixation categories reflect the distribution of objects likely to be of interest during daily tasks.
To estimate ET accuracy requirements, we computed the entire distribution of object visual sizes recorded in camera projection space. To place gaze on an object of average (projected) size 50% of the time, we measure at the distribution’s 50% mark. These measurements can help to inform system design; while 50% reliability may be useful supporting context in a broader contextual AI agent, a system relying heavily on ET may aim for higher coverage. The distributions for each interaction scenario are seen in Figure 3.
Recall that we use angular radius to represent the average case ET requirements, and minor axis span as a more conservative measurement. The accuracy of wearable ET devices is known to suffer in dynamic conditions (Onkhar et al., 2023), yet recent devices remain quite accurate in unconstrained settings [^5] https://www.tobii.com/products/eye-trackers/wearables/tobii-pro-glasses-3666https://pupil-labs.com/products/neon. Assuming a device with $\leq 3\degree$ accuracy during daily wear, our results indicate that the majority of fixated objects (radius average=$4.07\degree$; minor axis=$3.12\degree$), the majority of objects in the near-field (radius=$5.88\degree$; minor axis=$4.69\degree$), and nearly all interacted objects (radius=$10.81\degree$; minor axis=$9.10\degree$) are reliable for placing gaze on the correct object. Conversely, objects in the mid-field (radius=3.3°3.3°$\degree3$; minor axis=$2.54\degree$) will be somewhat unreliable at this signal quality, where roughly half of objects are not able to be detected.
4.2. Contextualized vision-language model queries
Figure 4. Experiments where prior gaze fixation contents are supplied to a VLM along with egocentric images. When many fixations given as context, the model can synthesize information between the image + gaze context to outperform a greedy baseline that only considers contents from the prompt. The error surfaces in light blue represent 95% confidence intervals.
Because we constrain the VLM model to respond by selecting from the list of all currently visible objects, these experiments are classification tasks where accuracy = correct selections / all trials. Cases where the VLM response failed to return parseable JSON ($<$1% of trials) were discarded.
We compare against a number of baselines to see if the model effectively uses the context supplied. These baselines implement simple heuristics on the image and gaze context. The lowest performing baseline is random guessing among all visible objects. We also implement random guesses from the list of previously fixated objects, and a greedy strategy which always chooses the most fixated object from a sequence of previously fixated objects. Note these baselines still use context provided from ET, but make no attempt to synthesize with the egocentric image for better world understanding.
4.2.1. E1: “What am I looking at?”
When querying the VLM only with the egocentric image as context, the model successfully predicts the current fixated object 10.3% (95% CI = [8.3%, 12.3%]) of the time (see Figure 4 (left)). While this surpasses random guessing from visible objects, it is quite low. An effective strategy was to always return the immediately preceding fixation (left tail of Random (prior fixations) in Figure 4). VLMs are not expected to excel at gaze prediction, as they are known to misinterpret context or hallucinate (Cui et al., 2023; Leng et al., 2024). However, including context from prior gaze greatly improves the model’s ability to predict the current fixation, with a peak accuracy of 24.8% (CI = [22.1%, 27.7%]) at 6 prior fixations. The gaze context-based baselines slightly outperform the VLM with one or few prior fixations, reinforcing that current gaze is contingent on scanpath history (Burlingham et al., 2024a, b). With more context (6+ fixations), the model begins to outperforms baselines, indicating that prior context and image contents are being synthesized, and the combination of contextual cues increase the model’s performance.
4.2.2. E2: “What am I going to interact with?”
E2 sees similar trends to E1; however, the more contextually-grounded task to predict the object of physical interaction sees greater benefit when including ET context. Clearly, prior eye gaze is a strong indicator for interaction, and by supplying historical gaze context, we can greatly improve the VLM’s ability to understand the user’s actions. We see a peak accuracy of 49.5% (CI = [43%, 56.1%]). As evident by this and the stronger baseline performances, gaze is tightly coupled with the onset of interaction. Note that we query only at positive examples where an interaction does take place, and the inclusion of a null case could have led to the model raising false positives / negatives.
@@ 5. Discussion
Our ET signal quality benchmark measures the likelihood of sensing objects, but has little considerations of edge cases (such as very small or very far objects). In the future, supplementary computations might aid the ability to place gaze on the correct object, possibly via contextual cues (Bi and Zhai, 2013) for error correction or additional sensors (Wei et al., 2023). These could alleviate ET sensor quality from becoming a bottleneck when using gaze to model human attention, specifically in more challenging contexts such as outdoors, where objects tend to be much further.

### v3

@@ 5. Discussion
Our ET signal quality benchmark measures the likelihood of sensing objects, but has little considerations of edge cases (such as very small or very far objects). In the future, supplementary computations might aid the ability to place gaze on the correct object, possibly via contextual cues (Bi and Zhai, 2013) for error correction or additional sensors (Wei et al., 2023). These could alleviate ET sensor quality from becoming a bottleneck when using gaze to infer human attention, specifically in more challenging contexts such as outdoors, where objects tend to be much further.
5.1. Future Work
While the ADT dataset provided a platform for simulating gaze-aided querying (Pan et al., 2023), it is critical to explore the impacts of eye tracking in real contextual AI scenarios. Prototypes leveraging point-in-time gaze have seen high user acceptance (Konrad et al., 2024; Wang et al., 2024), and the inclusion of temporal context is likely to better improve an agent’s ability to infer context and disambiguate a user’s queries. Thus, user experience investigations are important future avenues to follow up.
This work explored contextual inferences that could be made in short intervals lasting no longer than a few seconds, yet the information contained in larger time scales could also be invaluable for improving contextual AI agent understanding. Larger scales could enable better inferences of the current situation, and pave the way towards implicit personalization of contextual AI assistants (Pardini et al., 2022).